In New England, the horizontal component of the Earth's magnetic field has a magnitude of 1.2 × 10-5 T. An electron is shot vertically straight up from the ground with a speed of 3.4 × 106 m/s. What is the magnitude of the acceleration caused by the magnetic force? Ignore the gravitational force acting on the electron.
Answer #1

F=qvB

q=1.6×10-19c , V=3.4×106 , B=1.2×10-5 T

F=1.6×10-19×3.4×106×1.2×10-5

F=6.528×10-18

F=ma

a=F/m

a=(6.528×10-18)/9.1×10-31

a=7.173×1012 m/s^2

if you are satisfied from my solution then plz like it Thank you